# EigenLoRAx: Recycling Adapters to Find Principal Subspaces for Resource-Efficient Adaptation and Inference

### v2

@@ Abstract
The rapid growth of large models has raised concerns about their environmental impact and equity in accessibility due to significant computational costs. Low-Rank Adapters (LoRA) offer a lightweight solution for finetuning large models, resulting in an abundance of publicly available adapters tailored to diverse domains. We ask: Can these pretrained adapters be leveraged to further streamline adaptation to new tasks while addressing these challenges? We introduce EigenLoRAx, a parameter-efficient finetuning method that recycles existing adapters to create a principal subspace aligned with their shared domain knowledge which can be further augmented with orthogonal basis vectors in low-resource scenarios. This enables rapid adaptation to new tasks by learning only lightweight coefficients on the principal components of the subspace - eliminating the need to finetune entire adapters. EigenLoRAx requires significantly fewer parameters and memory, improving efficiency for both training and inference. Our method demonstrates strong performance across diverse domains and tasks, offering a scalable for edge-based applications, personalization, and equitable deployment of large models in resource-constrained environments.
@@ 1 Introduction
Recent advancements in natural language processing (Touvron et al., 2023) and computer vision (Rombach et al., 2021) have driven the rise of large-scale models with billions of parameters. However, the size and complexity of these models not only make it impractical for most researchers to train or fine-tune them on downstream tasks but also contribute significantly to their carbon footprint, raising concerns about environmental sustainability. To address these challenges, there has been growing interest in parameter-efficient finetuning (PEFT) methods, such as adapters (Houlsby et al., 2019; Chen et al., 2022; Luo et al., 2023), low rank adaptation (LoRA) methods (Hu et al., 2021; Kopiczko et al., 2023; Liu et al., 2024), prompt-based methods (Lester et al., 2021; Razdaibiedina et al., 2023; Fischer et al., 2024).
LoRA and its follow-up works (Meng et al., 2024; Liu et al., 2024) have gained significant attention for their simplicity. This has fueled the proliferation of thousands of low-rank adapters within the growing open-source community. Given that these adapters are underutilized, an important question arises: Can we recycle the information contained in them to improve the efficiency of subsequent tasks? Recent work has shown that weight updates in deep neural networks occur within low-dimensional invariant subspaces (Kwon et al., 2024), aligning with the universality hypothesis that neural network behavior and learned representations often reside in shared, structured subspaces (Chughtai et al., 2023; Guth & Ménard, 2024). This suggests that LoRA adapters may similarly share a principal subspace that can be reused, eliminating the need to rediscover it during the training of new adapters.
@@ 3 Method
In this section, we describe the theoretical motivation and the algorithm of our method, with a discussion on the hyper-parameters and quantification of practical benefits. Note that we use the terms EigenLoRA, EigenLoRAx, ELoRA and ELoRAx interchangeably.
@@ 3.1 Theoretical Motivation
Let $W\in\mathbb{R}^{m\times n}$ be a matrix representing a linear transformation between the vector spaces $\mathbb{R}^{m}$ and $\mathbb{R}^{n}$. When $W$ is of full rank, meaning $\text{rank}(W)=\min(m,n)$, it is capable of expressing the entire space of possible linear mappings between these dimensions. However, as the rank of $W$ increases, modifying it to accommodate new tasks becomes computationally expensive and increasingly complex. To mitigate this challenge in fine-tuning large models, LoRA introduces a decomposition of $W$ into two low-rank matrices, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$, where $r\ll\min(m,n).$ This factorization ensures that the product $BA$ retains the original dimensions of $W$ while having a significantly reduced rank. That is, $BA\in\mathbb{R}^{m\times n},\quad\text{with}\quad\text{rank}(BA)\leq r.$ As a result, although the transformation defined by $BA$ maps from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$, it does not span the full space of such mappings due to its constrained rank. By focusing on learning only the most essential updates within a low-dimensional subspace, LoRA provides a parameter-efficient mechanism for model adaptation. The number of parameters in this adaptation framework is given by $m\cdot r+r\cdot n$, which is substantially smaller than the full parameter count $m\cdot n$, making it a computationally viable alternative for fine-tuning large-scale models.
Many downstream adapters have been observed to converge toward a shared set of key directions in their parameter space (Meng et al., 2024; Liu et al., 2024), suggesting the existence of a principal subspace underlying diverse tasks. We hypothesize that LoRA adapters leverage such subspaces, enabling task-specific adaptations within a lower-dimensional structure rather than the full weight space. This not only reduces computational overhead but also reinforces the idea that task-relevant transformations reside within a compact, reusable subspace. To formalize this, we first define a space of tasks representable by linear transformation matrices, providing a foundation for analyzing the role of shared principal subspaces in model adaptation.
Definition 3.1 (Linear Transformation Tasks).
We define a task domain $\mathcal{T}_{d}$ as set of $d$ linear tasks, $\mathcal{T}_{d}=\{t\mid\mathcal{X}_{t}\in\mathbb{R}^{m}\rightarrow\mathcal{Y}_{t}\in\mathbb{R}^{n}\}$ where for each task $t\in T_{d}$, $\exists W^{*}_{t}\in\mathbb{R}^{m\times n}$ such that $Y=W^{*}_{t}X+\epsilon_{t}$ $\forall X\in\mathcal{X}_{t},Y\in\mathcal{Y}_{t}$ where $\epsilon_{t}$ denotes the noise in data distribution.
@@
We define the weights of a LoRA adapted for task $t$ as $B_{t}A_{t}$. Both $B_{t}$ and $A_{t}$ will have their own individual subspaces. For the purpose of the analysis we will consider a generic task specific weight matrix $W_{t}\in\mathbb{R}^{m\times n}$ adapted to task $t$ such that $n<m$ and its rank $r<n$. The analysis, however, is valid for both $B_{t}$ and $A_{t}$. Now can define a set of LoRAs as stacked (along columns) weight matrices $\hat{W}=\{W_{t}\}$ each adapted for a task $t\in\mathcal{T}_{d}$ and a training set $\mathcal{S}_{t}=\{\{x_{i},y_{i}\}\mid x_{i}\in\mathcal{X}_{t},y_{i}\in\mathcal{Y}_{t}\}^{s_{t}}_{i=1}$ where $s_{t}=|\mathcal{S}_{t}|$ and $\mathcal{X}_{t}\times\mathcal{Y}_{t}$ are distributed according to some unknown Gaussian distribution with mean $\bar{\mathcal{X}_{t}}$ and $\lVert X\lVert_{F}\leq M$ for some constant $M>0$. Each weight matrix can have different ranks and the following method and analysis will still hold, however, for brevity we assume all weight matrices stacked in $\hat{W}$ to have the same rank $r$.
@@
Definition 3.4 (Principal Subspace of LoRAs finetuned in domain $\mathcal{T}_{d}$).
We define the principal subspace of weights for a task domain $\mathcal{T}_{d}$ as $\mathcal{Z}^{K}_{d}=\{\alpha\mathcal{V}^{T}_{K}\mid\alpha\in\mathbb{R}^{m\times K}\}$ spanned by top K principal components of the LoRAs within $\mathbb{R}^{m\times n}$.
@@
. A new linear task $t_{d+1}$ with true solution $W^{*}$ is said to be related if it is spanned by the basis of $\hat{W}$ i.e. $W^{*}=C\hat{W}$ and it holds that $\lVert W^{*}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}\leq\|W^{*}-\alpha_{d+1}\mathcal{V}^{T}_{K}\lVert_{F}$ for all rank $K$ linear transformation matrices $\alpha_{d+1}$ and $\lVert W^{*}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}\leq\lVert C\lVert^{2}_{2}\sum_{i=K+1}^{nd}\sigma^{2}_{i}$ where $\sigma_{i}$’s are singular values of $\hat{W}$. For such a task, we learn coefficients of $K$ principal components $\alpha_{d+1}\in\mathbb{R}^{m\times K}$ resulting in EigenLoRAx weights $W_{\mathcal{E}}=\alpha_{d+1}\mathcal{V}^{T}_{K}$.
@@ 3.2 Algorithm
Assume we have $N$ LoRA adapters, each consisting of a set of $A,B$ matrix pairs for every layer, trained on various tasks within a domain $\mathcal{T}_{d}$ for a given base pretrained model. Algorithm 1 computes a list of top $K$ principal components—referred to as EigenLoRAx PCs—that define an initial principal subspace for this domain. To construct this subspace, the algorithm aggregates LoRA matrices across tasks for each layer, separately for $A$ and $B$ matrices (though it can also be applied to the product $BA$). Each LoRA matrix, having rank r, is treated as a list of vectors, and a decomposition is performed on this stacked set of vectors. The most significant components extracted from this process serve as a basis for the principal subspace, providing an efficient representation that can be linearly combined to approximate the original LoRA weight matrices. We illustrate our algorithm using generic weight matrices $W_{t}$, where each $W_{t}$represents a single $A$ or $B$ matrix from a single layer of the neural network. In practice, this procedure is applied to all relevant layers.
@@
$h=W_{0}x+\boxed{\alpha_{B}^{T}\mathcal{V}_{B}\alpha_{A}^{T}\mathcal{V}_{A}(x)}.$ (1)
Here, $W_{0}$ are the pre-trained base model weights and $\mathcal{V}_{B},\mathcal{V}_{A}$ are EigenLoRAx PCs that remain fixed during training. The corresponding $\alpha_{B}$ and $\alpha_{A}$ are learned. This reduces the number of learnable parameters from $O(rn)$ to $O(rK)$, by a factor of $\frac{n}{K}$ (assuming rank $r$ to be fixed, which could also be changed).
@@
For a task $t_{d+1}$, we assume a hypothesis $h\in\mathcal{H}_{W_{d+1}}$ expressed as $h(W_{d+1},X)=W_{d+1}X+W_{0}X+b$ where $W_{d+1}$ has rank $m$, $b$ is some constant and $W_{0}$ represents weights of a pretrained foundation model that is frozen during finetuning and $X,Y$ are sampled from $\mathcal{X}_{t},\mathcal{Y}_{t}$ respectively. We have $h^{\mathcal{E}}\in\mathcal{H}_{W_{\mathcal{E}}},h^{*}\in\mathcal{H}_{W}$ such that $h^{\mathcal{E}}(W_{\mathcal{E}},X)=\alpha_{d+1}\mathcal{V}^{T}_{K}X+W_{0}X+b$ where $W_{\mathcal{E}}$ has rank $K$ and $h^{*}(W^{*},X)=C\hat{W}X+W_{0}X+b$ where $h^{*}(W_{d+1},X)=Y$ is the true solution for task $t_{d+1}$. For a Lipschitz continuous loss ($\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h_{W})$) that is strong convex within the principal subspace defined by principal components $\mathcal{V}^{T}_{K}$ with Lipschitz constant ($L$), the risk can be written as $R_{\mathcal{S}_{t}}^{F}(h_{W})=E_{[}\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h_{W})]$ , and using Rademacher complexity bounds we can say with probability at least $1-4\delta$ for some $\delta>0$,
$\|W^{*}-W_{d+1}\lVert^{2}_{F}\leq C_{1}\cdot\Bigg{(}\frac{\sqrt{m}}{\sqrt{s_{t}}}\Bigg{)}+C_{2}$ (2)
@@
where $\sigma_{i}$ are singular values of $\hat{W}$, $C$ is some constant such that $W^{*}=C\hat{W}$ and $C_{1}$, $C_{2}$ are some constants.
@@
$R_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})-R_{\mathcal{S}_{t}}^{F}(h^{*})=\mathbb{E}_{\mathcal{S}_{t}}\big{[}\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})-\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h_{W}^{*})\big{]}$
@@
$\mathbb{E}_{\mathcal{S}_{t}}\big{[}\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})-\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h_{W}^{*})\big{]}\geq\frac{\mu}{2}\lVert W_{\mathcal{E}}-W^{*}\|^{2}_{F}$
@@
$|R_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})-\hat{R}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})|\leq\frac{\mathcal{R}_{s_{t}}(\mathcal{H}_{W_{\mathcal{E}}})}{2}+\sqrt{\frac{\ln(1/\delta)}{2s_{t}}}$
@@
$R_{\mathcal{S}_{t}}^{F}(h^{*})-R_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})=$ $R_{\mathcal{S}_{t}}^{F}(h^{*})-\hat{R}_{\mathcal{S}_{t}}^{F}(h^{*})-R_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})$
$+\hat{R}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})+\hat{R}_{\mathcal{S}_{t}}^{F}(h^{*})-\hat{R}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})$
Since we know by definition of $h^{*}$ that $\hat{R}_{\mathcal{S}_{t}}^{F}(h^{*})\leq\hat{R}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})$, we can say
$R_{\mathcal{S}_{t}}^{F}(h^{*})-R_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})\leq R_{\mathcal{S}_{t}}^{F}(h^{*})$ $-\hat{R}_{\mathcal{S}_{t}}^{F}(h^{*})$
$-R_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})+\hat{R}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})$
@@
$R_{\mathcal{S}_{t}}^{F}(h^{*})-R_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})\leq\mathcal{R}_{s_{t}}(\mathcal{H}_{W_{\mathcal{E}}})+\sqrt{\frac{2\ln(1/\delta)}{s_{t}}}$
@@
$\frac{\mu}{2}\lVert W^{*}-W_{\mathcal{E}}\lVert^{2}_{F}\leq\mathcal{R}_{s_{t}}(\mathcal{H}_{W_{\mathcal{E}}})+\sqrt{\frac{2\ln(1/\delta)}{s_{t}}}$ (4)
@@
$\mathcal{R}_{s_{t}}(\mathcal{H}_{W_{\mathcal{E}}})$ $=\mathcal{O}(\frac{\sqrt{K}\lVert W_{\mathcal{E}}\lVert_{F}}{\sqrt{s_{t}}})$
$=C_{1}\cdot\Bigg{(}\frac{\sqrt{K}\lVert W_{\mathcal{E}}\lVert_{F}}{\sqrt{s_{t}}}\Bigg{)}+C_{2}$
For a normalised $\lVert W_{\mathcal{E}}\lVert$ is usually bounded Hence,
$\lVert W^{*}-W_{\mathcal{E}}\lVert^{2}_{F}\leq C_{1}\cdot\Bigg{(}\frac{\sqrt{K}}{\sqrt{s_{t}}}\Bigg{)}+C_{2}$ (5)
@@
$\lVert W^{*}-W_{d+1}\lVert^{2}_{F}\leq C_{1}\cdot\Bigg{(}\frac{\sqrt{m}}{\sqrt{s_{t}}}\Bigg{)}+C_{2}$ (6)
This proves 2. Now to further prove 3, we use properties of Frobenius norm,
$\lVert W_{\mathcal{E}}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}-\lVert W^{*}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}\leq\|W_{\mathcal{E}}-W^{*}\lVert^{2}_{F}$
Then following from the definition of $W^{*}$, we can say that,
$\lVert W_{\mathcal{E}}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}-\lVert C\lVert^{2}_{2}\sum_{i=K+1}^{nd}\sigma^{2}_{i}\leq\|W_{\mathcal{E}}-W^{*}\lVert^{2}_{F}$
@@
$\lVert\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}$ $-W_{\mathcal{E}}\lVert^{2}_{F}\leq\lVert W^{*}-W_{\mathcal{E}}\lVert^{2}_{F}$
@@
We can just rewrite $W_{\mathcal{E}}=\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}$ and get the same bound as above for $\lVert\alpha^{*}_{d+1}-\alpha_{d+1}\lVert^{2}_{F}$. We can similarly obtain the upper bound for 3

### v3

@@ Abstract
The rapid growth of large models has raised concerns about their environmental impact and equity in accessibility due to significant computational costs. Low-Rank Adapters (LoRA) offer a lightweight solution for finetuning large models, resulting in an abundance of publicly available adapters tailored to diverse domains. We ask: Can these pretrained adapters be leveraged to further streamline adaptation to new tasks while addressing these challenges? We introduce EigenLoRAx, a parameter-efficient finetuning method that recycles existing adapters to create a principal subspace aligned with their shared domain knowledge which can be further augmented with orthogonal basis vectors in low-resource scenarios. This enables rapid adaptation to new tasks by learning only lightweight coefficients on the principal components of the subspace - eliminating the need to finetune entire adapters. EigenLoRAx requires significantly fewer parameters and memory, improving efficiency for both training and inference. Our method demonstrates strong performance across diverse domains and tasks, offering a scalable for edge-based applications, personalization, and equitable deployment of large models in resource-constrained environments. The code is available here.
@@ 1 Introduction
Recent advancements in machine learning have driven the rise of large-scale models with billions of parameters. However, the size and complexity of these models not only make it impractical for most researchers to train or fine-tune them on downstream tasks but also contribute significantly to their carbon footprint, raising concerns about environmental sustainability. To address these challenges, there has been growing interest in parameter-efficient finetuning (PEFT) methods, such as adapters (Houlsby et al., 2019; Chen et al., 2022; Luo et al., 2023), low rank adaptation (LoRA) methods (Hu et al., 2021; Kopiczko et al., 2023; Liu et al., 2024), prompt-based methods (Lester et al., 2021; Razdaibiedina et al., 2023; Fischer et al., 2024). LoRA and its follow-up works (Meng et al., 2024; Liu et al., 2024) have gained significant attention for their simplicity. This has fueled the proliferation of thousands of low-rank adapters within the growing open-source community. Given that these adapters are underutilized, an important question arises: Can we recycle the information contained in them to improve the efficiency of subsequent tasks? Recent work has shown that weight updates in deep neural networks occur within low-dimensional invariant subspaces (Kwon et al., 2024), aligning with the universality hypothesis that neural network behavior and learned representations often reside in shared, structured subspaces (Chughtai et al., 2023; Guth & Ménard, 2024). This suggests that LoRA adapters may similarly share a principal subspace that can be reused, eliminating the need to rediscover it during the training of new adapters.
@@ 3 Method
In this section, we present the theoretical foundation Section 3.1 and algorithmic details Section 3.2 of our method, followed by a discussion on hyperparameter selection and an assessment of its practical advantages. Note that we use the terms EigenLoRA, EigenLoRAx, ELoRA and ELoRAx interchangeably.
@@ 3.1 Theoretical Motivation
For a full rank weight matrix $W\in\mathbb{R}^{m\times n}$ that learns to map input space $X\in\mathbb{R}^{m}$ to output space $\mathbb{R}^{n}$, the rank is expressed as $\min(m,n)$. As the rank of $W$ increases, modifying it to accommodate new tasks becomes computationally expensive and increasingly complex. This is a common challenge faced when finetuning pretrained large foundation models. LoRA is a parameter efficient finetuning approach used for large pretrained models with weights $W_{0}$ that mitigates this challenge by merely learning low-rank weight updates $W$ such that the risk between $Y$ and $W_{0}X+WX+b$ is minimized. Instead of directly learning $W$, LoRA proposes to learn a lower ranked decomposition of $W$ by learning two low-rank matrices, $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$, both having ranks $r$. This factorization ensures that the product $BA$ retains the original dimensions of $W_{0}$ while having a significantly reduced rank. As a result, although the transformation defined by $BA$ maps from $\mathbb{R}^{m}$ to $\mathbb{R}^{n}$, it does not span the full space of such mappings due to its constrained rank. The low-rank weight matrices result in substantially smaller number of trainable parameters than the full rank parameter count of $m\cdot n$. Such parameter efficient finetuning makes LoRA a computationally viable alternative for fine-tuning large-scale models.
Previous works such as (Meng et al., 2024; Liu et al., 2024) have proposed the existence of a common parameter subspace implying the idea of shared principal subspace. We highlight that LoRA adapters share such a lower dimensional shared principal subspace when finetuned for diverse tasks. Along with reduction in computational overhead, it reinforces the idea that task-relevant transformations reside within a compact, reusable subspace. To formalize this, we first define a space of tasks representable by linear transformation matrices, providing a foundation for analyzing the role of shared principal subspaces in model adaptation.
Definition 3.1 (Task definition for LoRAs).
We first define a LoRA task $t_{i}(X_{i},Y_{i}):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ such that $Y_{i}=W^{*}_{i}X_{i}+b$ where $b$ is some constant. Then the LoRA task domain $\mathcal{T}_{d}$ is a set of $d$ such tasks, $\mathcal{T}_{d}=\{t_{i}\}^{d}_{i=1}$.
@@
We define the weights of a LoRA adapted for task $t_{i}$ as $B_{i}A_{i}$. Both $B_{i}$ and $A_{i}$ will have their own individual subspaces. For the purpose of the analysis we will consider a generic task specific weight matrix $W_{i}\in\mathbb{R}^{m\times n}$ adapted to task $t_{i}$ such that $n<m$ and its rank $r<n$. The analysis, however, is valid for both $B_{i}$ and $A_{i}$. Now can define a set of LoRAs as stacked (along columns) weight matrices $\hat{W}=\{W_{i}\}^{d}_{i=1}$ where each $W_{i}$ is adapted for a task $t_{i}\in\mathcal{T}_{d}$ and a training set $\mathcal{S}_{i}=\{\{x,y\}\mid x\in X_{t},y\in Y_{t}\}$ where the size of the training is $s_{i}=|\mathcal{S}_{i}|$. For theoretical analysis we assume that each training set $X_{i}\times Y_{i}$ is distributed according to some unknown Gaussian distribution with mean $\bar{X_{i}}$ and $\lVert X_{i}\lVert_{F}\leq M$ for some constant $M>0$. Each weight matrix can have different ranks and the following method and analysis will still hold, however, for brevity we assume all weight matrices stacked in $\hat{W}$ to have the same rank $r$.
@@
Definition 3.4 (Shared principal subspace of LoRAs finetuned in domain $\mathcal{T}_{d}$).
We define the shared principal subspace of weights for a task domain $\mathcal{T}_{d}$ as $\mathcal{Z}^{K}_{d}=\{\alpha\mathcal{V}^{T}_{K}\mid\alpha\in\mathbb{R}^{m\times K}\}$ spanned by top K principal components of the LoRAs within $\mathbb{R}^{m\times n}$.
@@
. A new linear task $t_{d+1}$ with true solution ${W^{*}}_{d+1}$ is said to be related if it is spanned by the basis of $\hat{W}$ i.e. ${W^{*}}_{d+1}=C\hat{W}$ and it holds that $\lVert{W^{*}}_{d+1}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}\leq\|{W^{*}}_{d+1}-\alpha_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}$ for all rank $K$ linear transformation matrices $\alpha_{d+1}$ and $\lVert{W^{*}}_{d+1}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}\leq\lVert C\lVert^{2}_{2}\sum_{i=K+1}^{nd}\sigma^{2}_{i}$ where $\sigma_{i}$’s are singular values of $\hat{W}$. For such a task, we learn coefficients of $K$ principal components $\alpha_{d+1}\in\mathbb{R}^{m\times K}$ resulting in EigenLoRAx weights $W_{\mathcal{E}}=\alpha_{d+1}\mathcal{V}^{T}_{K}$.
@@ 3.2 Algorithm
Assume we have $N$ LoRA adapters, each consisting of a set of $A,B$ matrix pairs for every layer, trained on various tasks within a domain $\mathcal{T}_{d}$ for a given base pretrained model. Algorithm 1 computes a list of top $K$ principal components—referred to as EigenLoRAx PCs—that define an initial principal subspace for this domain.
To construct this subspace, the algorithm aggregates LoRA matrices across tasks for each layer, separately for $A$ and $B$ matrices (though it can also be applied to the product $BA$). Each LoRA matrix, having rank r, is treated as a list of vectors, and a decomposition is performed on this stacked set of vectors. The most significant components extracted from this process serve as a basis for the principal subspace, providing an efficient representation that can be linearly combined to approximate the original LoRA weight matrices. We showcase our algorithm using representative weight matrices $W_{t}$, where each $W_{t}$ represents a single $A$ or $B$ matrix from a single LoRA layer of the neural network. In practice, this procedure is applied to all relevant layers.
@@
[…]
Here, $W_{0}$ are the pretrained weights of the base model and $\mathcal{V}_{B},\mathcal{V}_{A}$ are EigenLoRAx components (which represent the shared subspace) that are frozen during training. The corresponding lightweight coefficients $\alpha_{B}$ and $\alpha_{A}$ are learned. This reduces the number of learnable parameters from $O(2rn)$ to $O(2K)$, by a factor of $\frac{rn}{K}$ (assuming $\alpha$ to be scalar).
@@
For a task $t_{d+1}$, we assume a hypothesis $h\in\mathcal{H}_{W_{d+1}}$ expressed as $h(W_{d+1},X)=W_{d+1}X_{d+1}+W_{0}X_{d+1}+b$ where $W_{d+1}$ has rank $m$, $b$ is some constant and $W_{0}$ represents weights of a pretrained foundation model that is frozen during finetuning respectively. We have $h^{\mathcal{E}}\in\mathcal{H}_{W_{\mathcal{E}}},h^{*}\in\mathcal{H}_{{W^{*}}_{d+1}}$ such that $h^{\mathcal{E}}(W_{\mathcal{E}},X_{d+1})=\alpha_{d+1}\mathcal{V}^{T}_{K}X_{d+1}+W_{0}X_{d+1}+b$ where $W_{\mathcal{E}}$ has rank $K$ and $h^{*}({W^{*}}_{d+1},X_{d+1})=C\hat{W}X_{d+1}+W_{0}X_{d+1}+b$ where $h^{*}({W^{*}}_{d+1},X_{d+1})=Y_{d+1}$ is the true solution for task $t_{d+1}$. For a Lipschitz continuous loss ($\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h)$) that is strong convex within the shared principal subspace spanned by principal components $\mathcal{V}^{T}_{K}$ with some Lipschitz constant ($L$), the risk can be written as $R_{\mathcal{S}_{d+1}}^{F}(h_{W})=E_{\mathcal{S}_{t}}[\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h)]$ , and using Rademacher complexity bounds we can say with probability at least $1-4\delta$ for some $\delta>0$,
$\|{W^{*}}_{d+1}-W_{d+1}\lVert^{2}_{F}\leq C_{1}\cdot\Bigg{(}\frac{\sqrt{m}}{\sqrt{s_{t}}}\Bigg{)}+C_{2}$ (1)
@@
where $\sigma_{i}$ are singular values of $\hat{W}$, $C$ is some constant such that ${W^{*}}_{d+1}=C\hat{W}$ and $C_{1}$, $C_{2}$ are some constants.
@@
$R_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})-R_{\mathcal{S}_{d+1}}^{F}(h^{*})=\mathbb{E}_{\mathcal{S}_{t}}\big{[}\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})-\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h^{*})\big{]}$
@@
$\mathbb{E}_{\mathcal{S}_{t}}\big{[}\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h^{\mathcal{E}})-\mathcal{\ell}_{\mathcal{S}_{t}}^{F}(h^{*})\big{]}\geq\frac{\mu}{2}\lVert W_{\mathcal{E}}-{W^{*}}_{d+1}\|^{2}_{F}$
@@
$|R_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})-\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})|\leq\frac{\mathcal{R}_{s_{d+1}}(\mathcal{H}_{W_{\mathcal{E}}})}{2}+\sqrt{\frac{\ln(1/\delta)}{2s_{t}}}$
@@
$R_{\mathcal{S}_{d+1}}^{F}(h^{*})-R_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})=$ $R_{\mathcal{S}_{d+1}}^{F}(h^{*})-\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{*})$
$-R_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})+\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})$
$+\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{*})-\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})$
Since we know by definition of $h^{*}$ that $\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{*})\leq\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})$, we can say
$R_{\mathcal{S}_{d+1}}^{F}(h^{*})-R_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})\leq R_{\mathcal{S}_{d+1}}^{F}(h^{*})-\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{*})$
$-R_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})+\hat{R}_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})$
@@
$R_{\mathcal{S}_{d+1}}^{F}(h^{*})-R_{\mathcal{S}_{d+1}}^{F}(h^{\mathcal{E}})\leq\frac{\mathcal{R}_{s_{d+1}}(\mathcal{H}_{W_{\mathcal{E}}})}{2}+\sqrt{\frac{2\ln(1/\delta)}{s_{d+1}}}$
$+\frac{\mathcal{R}_{s_{d+1}}(\mathcal{H}_{{W^{*}}_{d+1}})}{2}$
@@
$\frac{\mu}{2}\lVert{W^{*}}_{d+1}-W_{\mathcal{E}}\lVert^{2}_{F}\leq\frac{\mathcal{R}_{s_{d+1}}(\mathcal{H}_{W_{\mathcal{E}}})}{2}+\sqrt{\frac{2\ln(1/\delta)}{s_{t}}}+\frac{\mathcal{R}_{s_{d+1}}(\mathcal{H}_{{W^{*}}_{d+1}})}{2}$ (3)
@@
$\mathcal{R}_{s_{d+1}}(\mathcal{H}_{W_{\mathcal{E}}})$ $=\mathcal{O}(\frac{\sqrt{K}\lVert W_{\mathcal{E}}\lVert_{F}}{\sqrt{s_{t}}})$
We can separate the constants including $\mathcal{R}_{s_{d+1}}(\mathcal{H}_{{W^{*}}_{d+1}})$ from 3 and assume that, for a normalised $\lVert W_{\mathcal{E}}\lVert$, it is usually bounded, then we can write:
$\lVert{W^{*}}_{d+1}-W_{\mathcal{E}}\lVert^{2}_{F}\leq C_{1}\cdot\Bigg{(}\frac{\sqrt{K}}{\sqrt{s_{t}}}\Bigg{)}+C_{2}$ (4)
@@
$\lVert{W^{*}}_{d+1}-W_{d+1}\lVert^{2}_{F}\leq C_{1}\cdot\Bigg{(}\frac{\sqrt{m}}{\sqrt{s_{t}}}\Bigg{)}+C_{2}$ (5)
This proves 1. Now to further prove 2, we use properties of Frobenius norm,
$\lVert W_{\mathcal{E}}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}-\lVert{W^{*}}_{d+1}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}\leq\|W_{\mathcal{E}}-{W^{*}}_{d+1}\lVert^{2}_{F}$
Then following from the definition of ${W^{*}}_{d+1}$, we can say that,
$\lVert W_{\mathcal{E}}-\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}\lVert^{2}_{F}-\lVert C\lVert^{2}_{2}\sum_{i=K+1}^{nd}\sigma^{2}_{i}\leq\|W_{\mathcal{E}}-{W^{*}}_{d+1}\lVert^{2}_{F}$
@@
$\lVert\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}$ $-W_{\mathcal{E}}\lVert^{2}_{F}\leq\lVert{W^{*}}_{d+1}-W_{\mathcal{E}}\lVert^{2}_{F}$
@@
We can just rewrite $W_{\mathcal{E}}=\alpha^{*}_{d+1}\mathcal{V}^{T}_{K}$ and get the same bound as above for $\lVert\alpha^{*}_{d+1}-\alpha_{d+1}\lVert^{2}_{F}$. We can similarly obtain the upper bound for 2